# BSMamba: Brightness and Semantic Modeling for Long-Range Interaction in Low-Light Image Enhancement

### v1

@@ 4 Experiments
Figure 10: Validation on Downstream Applications. We validate DE-Net’s effectiveness through edge detection, which is sensitive to image details, and Semantic Mamba’s effectiveness through instance segmentation. Observations show that our full model achieves the best performance in both edge detection and instance segmentation.
4.1 Datasets and Experimental Setting
Our method is evaluated on four widely recognized datasets for LLIE: LOL [47], LOLv2-Real [47], LSRW-Huawei [15], LSRW-Nikon [15]. The LOL dataset comprises 485 paired images for training and 15 pairs for testing, with LOLv2-Real [47] used for evaluation. The LSRW-Huawei dataset, captured with a Huawei P40 Pro, includes 3,150 training pairs and 20 test pairs, while the LSRW-Nikon dataset, collected using the Nikon D7500, contains 2,450 training pairs and 30 test pairs. Additionally, we evaluate on two unpaired datasets:DICM [22] (64 images) and NPE [37] (85 images).
The model is implemented and trained end-to-end using the PyTorch framework. During training, images are randomly cropped to $256\times 256$ pixels and augmented with random horizontal and vertical flips. We optimize the model using the ADAM optimizer, starting with a learning rate of $4.0\times 10^{-4}$, and employ a multi-step learning rate scheduler. The batch size is set to 8, and the training process spans $1.5\times 10^{5}$ iterations. All experiments are conducted on two NVIDIA 4090 GPUs.
4.2 Overall Performance
We compare our approach against a variety of state-of-the-art LLIE methods, including traditional algorithms (LIME [14], NPE [37], SRIE [8]), deep learning-based techniques (Kind [52], Kind++ [53], MIRNet [48], SGM [47], HDMNet [28], SNR-Aware [45], SNR-SKF [43], Retinexformer [2], UHDFormer [36]), Fourier-based models (FourLLIE [35], FECNet [18], UHDFour [24], DMFourLLIE [50]), and Mamba-based approaches (RetinexMamba [1], Wave-Mamba [55], MambaIR [13]). For fair evaluation, all deep learning models are trained on the same datasets using their respective publicly available implementations.
We adopt Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [39], and Learned Perceptual Image Patch Similarity (LPIPS) [49] as evaluation metrics. Higher PSNR and SSIM values, along with lower LPIPS scores, indicate better visual quality and closer resemblance to the ground truth.
@@ 4.2 Overall Performance
Quantitative Results on LOL, LOLv2-Real, LSRW-Huawei and LSRW-Nikon. 1) Tab. 1 reports the quantitative results on the LSRW-Nikon dataset, where our method achieves the best PSNR, SSIM, and LPIPS scores, significantly outperforming existing approaches. Compared to methods like RetinexMamba, our approach not only ensures a lightweight design but also substantially reduces computational overhead. 2) Tab. 2 presents results on the LSRW-Huawei dataset, where our method also achieves state-of-the-art (SOTA) performance, excelling in brightness restoration, structural preservation, and perceptual quality. Notably, compared to lightweight models such as UHDFormer and DMFourLLIE, our approach consistently delivers superior performance, highlighting its robustness and efficiency. 3) Tab. 3 shows the comparison results on the LOL-v1 and LOLv2-Real datasets. Our method achieves the highest PSNR, SSIM, and LPIPS scores, outperforming all competing methods.
@@
Qualitative Comparison. To intuitively demonstrate restoration quality, we provide visual comparisons with state-of-the-art methods. 1) Fig. 5 shows visual results on the LSRW-Nikon dataset. In local regions, existing methods exhibit noticeable blurring and color distortion, particularly in areas with fine textures. For example, methods like MambaIR and UHDFour fail to recover details and introduce artifacts. In contrast, our approach preserves color and luminance fidelity while maintaining clear texture structures, achieving high consistency with the Ground Truth. 2) Fig. 6 presents visual comparisons on the LSRW-Huawei dataset. For challenging local regions, our method outperforms others in restoring text details and object textures. Additionally, the color distribution is more natural, particularly in difficult areas. 3) Fig. 7 illustrates visual comparisons on the LOL dataset. In the first row, other methods misinterpret bright light sources, causing noticeable color shifts. In contrast, our approach preserves the true brightness of light sources while effectively enhancing darker areas, ensuring natural luminance balance and mitigating artifacts that degrade visual quality.
Qualitative Comparison on DICM and NPE Datasets. As shown in Fig.8, our method delivers superior visual results. In the first two rows, methods like UHDFormer and WaveMamba suffer from overexposure in brighter regions, while SNR-SKF introduces noticeable color distortions. In the third row, other methods cause significant texture blurring, whereas our approach preserves fine textures, enhances color fidelity, and improves illumination contrast. Additionally, our method demonstrates greater robustness and adaptability across diverse lighting conditions.
4.3 Ablation Study and Experimental Configurations
@@
We conduct ablation experiments to validate the independent contributions of each module in our framework. As shown in Fig.9, w/o represents the removal of specific modules: w/o Brightness Mamba significantly reduces brightness enhancement; w/o Semantic Mamba degrades semantic consistency and instance-level performance; w/o DE-Net results in blurred details and reduced edge sharpness. BSMamba Block $\rightarrow$ Vanilla SS2D replaces our Brightness and Semantic Mamba scanning strategies with a vanilla state-space 2D (SS2D) approach, causing severe degradation in both brightness and image details.
The performance of different configurations is presented in Tab.4. We evaluate the structural design of the Brightness Mamba and Semantic Mamba components: Brightness Mamba + Semantic Mamba modifies them to operate in parallel with summed outputs; Brightness Mamba C Semantic Mamba concatenates their outputs; Semantic Mamba $\leftrightarrow$ Brightness Mamba swaps their order. We also assess the impact of different brightness scoring mechanisms in the Brightness Mamba module: DA-CLIP $\rightarrow$ Brightness Histogram replaces DA-CLIP with a brightness histogram, calculated by normalizing pixel intensity distributions and using histogram bins as brightness scores; DA-CLIP $\rightarrow$ Y component uses the Y component from the YCrCb color space as the brightness score. Our full model, integrating all components and using DA-CLIP for brightness scoring, achieves the best performance, as shown in the last row.
4.4 Downstream Applications
@@
Beyond the brightness-focused Brightness Mamba, we introduce Semantic Mamba and DE-Net to enhance semantic consistency and detail restoration. To validate their effectiveness, we evaluate our method on downstream tasks, including instance segmentation (Mask R-CNN [17] pre-trained on COCO dataset [29]) and edge detection [4], which assess semantic preservation and detail recovery, respectively. Fig.9 illustrates the results of removing the corresponding modules and compares them. The visual comparisons further confirm the performance gains achieved by incorporating Semantic Mamba and DE-Net.
5 Conclusion
@@
In this paper, we propose BSMamba, a novel visual Mamba architecture designed to overcome the fundamental limitations of existing LLIE methods. By introducing the Brightness Mamba and Semantic Mamba, our approach redefines the label interaction paradigm based on brightness and semantic similarity, enabling effective long-range modeling of causally related labels while maintaining spatial and contextual consistency. Furthermore, BSMamba provides new insights and inspiration for future research. In the future, we plan to explore more generalized scanning strategies for brightness and semantic Mamba, or develop task-specific customized scanning strategies tailored to specific downstream tasks, aiming to guide and enhance the performance of related tasks.

### v2

@@ Introduction
Figure 9: Validation on Downstream Applications. We validate DE-Net’s effectiveness through edge detection, which is sensitive to image details, and Semantic Mamba’s effectiveness through instance segmentation.
Figure 10: Visual comparisons on unpair datasets.
Datasets and Experimental Setting
Our method is evaluated on four widely recognized datasets for LLIE: LOL (Yang et al. 2021), LOLv2-Real, LSRW-Huawei (Hai et al. 2023), and LSRW-Nikon, comprising a total of 485 training pairs and 15 test pairs for LOL, 3,150 training pairs and 20 test pairs for LSRW-Huawei, and 2,450 training pairs and 30 test pairs for LSRW-Nikon. Additionally, we assess performance on three unpaired datasets: NPE (Wang et al. 2013), MEF (Ma, Zeng, and Wang 2015), and VV.
The model is implemented and trained end-to-end using the PyTorch framework. During training, images are randomly cropped to $256\times 256$ pixels and augmented with random flips. We optimize the model using the ADAM optimizer with an initial learning rate of $4.0\times 10^{-4}$ and a multi-step learning rate scheduler. The batch size is set to $8$, and training spans $1.5\times 10^{5}$ iterations, conducted on two NVIDIA 4090 GPUs.
Overall Performance
We compare our approach against various state-of-the-art LLIE methods, including learning-based techniques (Kind (Zhang, Zhang, and Guo 2019), Kind++ (Zhang et al. 2021), MIRNet (Zamir et al. 2020), SGM (Yang et al. 2021), HDMNet (Liang et al. 2022), SNR-Aware (Xu et al. 2022), SNR-SKF (Wu et al. 2023), Retinexformer (Cai et al. 2023), UHDFormer (Wang et al. 2024a), CIDNet (Yan et al. 2025)), Fourier-based models (FourLLIE (Wang, Wu, and Zhi 2023), FECNet (Huang et al. 2022), UHDFour (Li et al. 2023), DMFourLLIE (Zhang et al. 2024a)), and Mamba-based approaches (RetinexMamba (Bai, Yin, and He 2024), Wave-Mamba (Zou et al. 2024a), MambaIR (Guo et al. 2025), MambaLLIE (Weng et al. 2025), CWNet (Zhang et al. 2025)). For fair evaluation, all deep learning models are trained on the same datasets using their publicly available implementations. We use Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) (Wang et al. 2004), and Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al. 2018) as evaluation metrics.
@@
$\lambda_{1},\lambda_{2},\lambda_{3}=[1.0,0.3,0.1]$ 21.51 0.6408 0.1637
$\lambda_{1},\lambda_{2},\lambda_{3}=[1.0,0.5,0.2]$ 21.29 0.6431 0.1688
@@
Quantitative Results on LOL, LOLv2-Real, LSRW-Huawei and LSRW-Nikon. Our method demonstrates advantages in model size and efficiency. 1) Tab. 1 presents results on the LSRW-Nikon dataset, where our approach significantly outperforms existing methods. 2) In Tab. 2, we achieved competitive results on the LSRW-Huawei dataset, with our PSNR ranking second, just 0.01 dB behind CWNet. 3) Tab. 3 shows that we also perform competitively on the LOL-v1 and LOL-v2-Real datasets.
@@
Qualitative Comparison. To illustrate restoration quality, we provide visual comparisons with state-of-the-art methods. 1) Fig.5 shows results on the LSRW-Nikon dataset, where existing methods suffer from blurring and color distortion, especially in fine textures. In contrast, our approach maintains color fidelity and clear texture details, aligning closely with the Ground Truth. 2) Fig.6 presents comparisons on the LSRW-Huawei dataset. Our method excels in restoring text and object textures in challenging regions, with a more natural color distribution. 3) Fig. 7 illustrates results on the LOL dataset. Other methods misinterpret bright light sources, leading to color shifts, while our approach preserves true brightness and effectively enhances dark areas, ensuring a natural luminance balance and reducing artifacts.
Ablation Study and Experimental Configuration
@@
Visual Ablation: As shown in Fig. 8, removing Brightness Mamba significantly compromises brightness enhancement. Excluding Semantic Mamba reduces semantic consistency and instance-level performance, while removing DE-Net leads to blurred details and decreased edge sharpness. Replacing our Brightness and Semantic Mamba scanning strategies with Vanilla SS2D (fixed multi-directional scanning) in the BSMamba Block results in severe declines in both brightness and image details.
Quantitative Ablation: Tab. 4 presents our ablation studies. The first set modifies and replaces core modules; the second set tests combinations of the two Mamba modules (addition, concatenation, reordering) to validate our optimal strategy. The third set assesses the impact of changing DA-CLIP’s brightness score to histogram and YCrCb luminosity statistics. The fourth group examines the effects of varying loss weights. These four groups comprehensively validate the effectiveness of our components and configurations.
@@
Downstream Applications: Beyond the brightness-focused Brightness Mamba, we introduce Semantic Mamba and DE-Net to enhance semantic consistency and detail restoration. To validate their effectiveness, we evaluate our method on downstream tasks, including instance segmentation (Mask R-CNN (He et al. 2017) pre-trained on COCO dataset (Lin et al. 2014)) and edge detection (Ding and Goshtasby 2001), which assess semantic preservation and detail recovery, respectively. Fig.8 illustrates the results of removing the corresponding modules and compares them. The visual comparisons further confirm the performance gains achieved by incorporating Semantic Mamba and DE-Net.
Limitations and Conclusion
Limitations: As observed in Fig. 10, our visual comparisons with current SOTA Mamba-based methods on paired datasets show that while we achieve favorable color and detail in the first two rows, we still encounter exposure issues in the third row. This insight motivates us to address the challenge of suppressing exposure artifacts while ensuring color semantic and brightness recovery in future work.
@@
Conclusion: In this paper, we introduced BSMamba, a novel visual Mamba architecture aimed at overcoming the fundamental limitations of LLIE. By incorporating brightness and semantic Mamba, our approach redefines the label interaction paradigm based on brightness and semantic similarity, enabling effective long-range modeling of causally related labels. In the future, we plan to explore generalized scanning strategies or develop customized scanning techniques tailored for specific downstream tasks to enhance performance.